OpenStudy (anonymous):

Which equation is an identity A. -3y+5=-3y-5 B. 7m-5=8m+7-m C. 15-(3v+7)=-3v-8 D. 5w+8-w=6-2(w-4)

OpenStudy (precal):

A. No solution 5=-5 B. No solution -5=7 C. No solution 8=-8 D. 4w+8=10-2w 6w=8 w=8/6 Are you sure you copied the question correctly? Identity means all real numbers